OpenStudy (anonymous):

Write the solution to the following problem as an ordered pair: The sum of two numbers is 170 and their difference is 26.

OpenStudy (xishem):

\[a+b=170\]\[a-b=26\]\[2a=196 \rightarrow a=98\]\[98+b=170 \rightarrow b=72\]

OpenStudy (xishem):

\[(98, 72)\]